OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6.

OpenStudy (anonymous):

Point slope equation: y-y1=m(x-x1) Then you plug in the numbers. slope=m, y1=the y given, x1= the x given y-(-1)=6(x-5), so y+1=6(x-5) and there you go!

OpenStudy (anonymous):

wait waht?

OpenStudy (anonymous):

The answer is y+1=6x-5

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

********* y+1=6(x-5)

OpenStudy (anonymous):

^^^^^ thats the answer

OpenStudy (anonymous):

thanks againn:)

OpenStudy (anonymous):

Your welcome!!